OpenStudy (anonymous):

According to this property, 3(r + 4) = 3r + 12 Answer (A) Commutative (B) Distributive (C) Multiplicative Identity (D) Multiplicative of Zero (E) Associative

OpenStudy (anonymous):

What do you think tell me first???

OpenStudy (anonymous):

Can you tell or not?? What do you think @LAGIRL69

OpenStudy (anonymous):

Commutative property: When two numbers are multiplied together, the product is the same regardless of the order of the multiplicands. For example 4 * 2 = 2 * 4 Associative Property: When three or more numbers are multiplied, the product is the same regardless of the grouping of the factors. For example (2 * 3) * 4 = 2 * (3 * 4) Multiplicative Identity Property: The product of any number and one is that number. For example 5 * 1 = 5. Distributive property: The sum of two numbers times a third number is equal to the sum of each addend times the third number. For example 4 * (6 + 3) = 4*6 + 4*3

OpenStudy (anonymous):

Distributive property says: \[\large a(b+c) = ab + ac\] Now what is your answer tell me..

OpenStudy (anonymous):

@LAGIRL69 it starts with B

OpenStudy (anonymous):

I think it is B @waterineyes

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

Then you are right..

OpenStudy (anonymous):

@laxplayer & @waterineyes Thanks Guys :)

OpenStudy (anonymous):

Welcome dear..

OpenStudy (anonymous):

No problem any time